The quality grade point average gpa of 200level economics student follows approximately a normal distribution with mean of 3.4 and SE of 2.2.How many of these students would expect to have a score between 3.5 and 4.5 inclusive,if the grade point average are computed to the nearest tenth

1
Expert's answer
2022-03-17T07:39:47-0400

"\mu=3.4, \sigma=2.2, n=200,"

"P(3.5≤x≤4.5)=P(\frac{3.5-3.4}{2.2}≤Z\leq \frac{4.5-3.4}{2.2})"

"=P(0.05\u2264Z\u22640.5)=P(Z\u22640.5)-P(Z\u22640.05)"

"=0.6915-0.5160=0.1755" ,

"0.1755*200=35.1"